OpenStudy (anonymous):

Solve using the elimination method show your work -x+4y=0.5 -3x-6y=10.5

OpenStudy (anonymous):

Okay i'd say it would be easiest if we divided every term in the second equation by -3 so now we have x + 2y = -3.5 Lets set them up to eliminate -x + 4y = 0.5 x + 2y = -3.5 x is eliminated 6y = -3 divide by 6 y = -3/6 or y = -1/2 now lets plug it in to solve for x -x + 4 (-1/2) = 0.5 -x -2 = 0.5 add 2 -x = 1.5 divide by -1 x = -1.5 so our solution is (-1.5, -0.5)

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

no problem.